(W+2)/(2W-2)=5 can you help solve for W in the steps so I can learn?
Answer by midwood_trail(310)

(W+2)/(2W-2)= 5 Can you help solve for W in steps so I can learn?
What we have here is called a fractional equation.
It's called fractional because the left side is a fraction. See it?
To remove that ugly fraction, we multiply both sides of the equation by the denominator, which is (2W - 2).
(2W-2) times [(W+2)/(2W-2)] = 5 times (2W-2)
Proceed the same way you would handle regular fractions.
On the left side, we cancel out anything that looks the same.
Do you see what I mean?
The quantity (2W-2) cancels out on the left side leaving us with
the quantity (W+2).
On the right side, we use the distributive rule.
Do you know what this means?
On the right side, we multiply 5 by EVERY term inside the quantity (2W-2).
Step by step:
5(2W-2) =
5 times 2W = 10W
5 times -2 = -10
On the right side, we now have this: 10W - 10.
Here is what your original equation has become:
W + 2 = 10W - 10
Are you with me so far?
Notice that your fractional equation has now become a linear equation.
We now simply solve for x. Don't forget to combine like terms!
W - 10W = -2 - 10
-9W = -12
To find the value of W, we now divide both sides of the equation by -9.
W = -12/-9
W = 12/9
Of course, your teacher may want you to reduce 12/9 to the lowest term.
In that case, 12/9 is reduced to 4/3.
Final answer: W = 4/3
Did you follow?